# Small Neutrino Masses from a Decoupled Singlet Scalar Field

### v1

@@ Abstract
We propose a unified solution with $Z_{4}$ discrete symmetry for small neutrino masses and stability of dark matter. Our model consists of an inert doublet scalar, a dark singlet scalar $S$, a spurion scalar $\varphi$ and right-handed neutrinos beyond the Standard Model (SM), which transform nontrivially under $Z_{4}$. Once the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of $\varphi$, much below the mass scale of the singlet scalar $S$, a small quartic coupling between SM Higgs and inert doublets, which violates both $Z_{4}$ and lepton number, is generated at tree level, so small neutrino masses are obtained at one-loops. We discuss the consistent parameter space for neutrino masses, direct detection and thermal leptogenesis in the model.
@@
Introduction: Neutrino oscillations ParticleDataGroup:2022pth indicate that neutrinos have nonzero masses about $m_{\nu}\sim 0.05\,{\rm eV}$. Moreover, from neutrinoless double beta decay experiments for Majorana neutrinos KATRIN:2021uub as well as Cosmic Microwave Background anisotropies and Baryon Acoustic Oscillations at cosmological scales Planck:2018vyg , we have set upper bounds on the neutrino masses as $m_{{\bar{\nu}}_{e}}<0.8\,{\rm eV}$ and $\sum_{i}m_{\nu_{i}}<0.12\,{\rm eV}$ at 95% CL, respectively. In order to account for small neutrino masses, we need to extend the Standard Model (SM) by the dimension-5 operators for Majorana neutrino masses with a cutoff scale much larger than the electroweak scale or generate neutrino masses of Dirac type from the Yukawa couplings with light sterile neutrinos much below the electroweak scale Cai:2017jrq .
@@
Seesaw mechanism Minkowski:1977sc ; Yanagida:1979as ; Gell-Mann:1979vob ; Mohapatra:1979ia gives rise to small neutrino masses due to the mixing between active neutrinos and heavy Majorana neutrinos. Alternatively, radiative seesaw scenarios provide small neutrino masses by loop corrections with new particles not far from the electroweak scale, due to small loop suppression factors. In the scotogenic model Ma:2006km , neutrino masses are generated by one-loop corrections with right-handed (RH) neutrinos and an extra doublet scalar, which can be a dark matter candidate Barbieri:2006dq ; LopezHonorez:2006gr . However, there is still a need of very small parameters in obtaining small neutrino masses such as a small mixing quartic coupling in the Higgs sector or small Yukawa couplings for the RH neutrinos.
@@
In this article, we propose a novel model with $Z_{4}$ discrete Peccei-Quinn (PQ) symmetry for the common origin of small neutrino masses and stability of dark matter. In our model, there are one inert doublet scalar, three right-handed neutrinos, a real spurion scalar and a singlet complex scalar field $S$, all transforming nontrivially under $Z_{4}$. The lepton number is respected in the renormalizable interactions due to $Z_{4}$. Then, after the $Z_{4}$ symmetry is broken to $Z_{2}$ due to the VEV of the spurion field, the lepton number is broken, and the lightest neutral dark scalar in the model becomes a dark matter candidate.
@@
Taking the decoupling limit for making the singlet scalar field $S$, we show how a small mass splitting between the neutral components of the inert doublet arises, in turn giving rise to small neutrino masses at one-loops. Our results remain intact for any higher order discrete symmetries or a continuous approximate $U(1)$ PQ symmetry containing $Z_{2}$ as a subgroup or in the cases with more complex singlet scalars and inert doublet scalars. We discuss various phenomenological constraints on the $Z_{4}$ violation from neutrino masses, dark matter direction detection and thermal leptogenesis.
@@
Discrete symmetry and dark matter scalars: We introduce a complex singlet scalar $S$ and one extra electroweak doublet scalar, $H_{2}$, other than the SM Higgs doublet $H_{1}$, and three right-handed (RH) neutrinos, $N_{Ri}$, with $i=1,2,3$, and a spurion real scalar field $\varphi$, beyond the SM. We impose a $Z_{4}$ PQ symmetry for the extra fields such that $\phi_{i}\to e^{iQ_{X}\pi/2}\phi_{i}$ with $\phi_{i}=S,\varphi,H_{1},H_{2},N_{R}$ and $Q_{X}(\phi_{i})=(+1,+2,0,+1,+1)$. On the other hand, the SM fermions are neutral under $Z_{4}$ so the model is different from the flavor-dependent $Z_{4}$ symmetry introduced in Ref. Babu:2013pma . As the subgroup of the $Z_{4}$ symmetry, we can identify $Z_{2}$ parities as $Z_{2}(\phi_{i})=(-1,+1,+1,-1,-1)$. The spurion field $\varphi$ transforms by $\varphi\to-\varphi$ under the $Z_{4}$ symmetry but it is invariant under the $Z_{2}$ subgroup. The spurion field $\varphi$ can be a fundamental scalar field or a composite field due to hidden strong dynamics.
@@
with ${\widetilde{H}}_{2}=i\sigma_{2}H^{*}_{2}$. It is remarkable that $\frac{1}{2}\lambda_{5}(H^{\dagger}_{1}H_{2})^{2}$ and its complex conjugate, vanish due to the $Z_{4}$ symmetry, and there is no bare mass terms for right-handed neutrinos. We note that there are extra $Z_{4}$ invariant interaction terms in the potential such as $\mu S^{2},\frac{1}{2}\lambda_{H_{i}\varphi}\varphi^{2}|H_{i}|^{2}$($i=1,2$), etc, but it is sufficient to focus on eq. (1) for our discussion below.
@@
where $\kappa^{\prime}\equiv\lambda^{\prime}_{S\varphi}v_{\varphi}$ is the $Z_{4}$ breaking effective interaction and $M_{N,i}\equiv\lambda_{N,ii}v_{\varphi}$ are the masses for RH neutrinos. Here, we assumed the basis where $\lambda_{N,ij}$ is diagonalized. After electroweak symmetry breaking with $\langle H_{1}\rangle=(0,v_{H})^{T}/\sqrt{2}\neq 0$, there appear mass mixings between dark matter scalars, namely, $S$ and the neutral components of $H_{2}$, so nonzero neutrino masses can be generated at one-loops, which will be discussed later.
We remark that $S$ and $H_{2}$ do not get VEVs, leaving a $Z_{2}$ symmetry unbroken. $S$, $H_{2}$ and $N_{Ri}$ are odd under the remaining $Z_{2}$, which ensures the stability of the lightest neutral $Z_{2}$-odd particle among them as a dark matter candidate. In the following discussion, in order to explain small neutrino masses without introducing small parameters, we consider the case where the singlet scalar $S$ is decoupled. In this case, either the lighter neutral component of $H_{2}$ or the lightest RH neutrino can be a dark matter candidate.
Figure 1: Feynman diagrams contributing to the effective $\lambda_{5,{\rm eff}}$ at tree level.
Decoupled scalar and inert doublet scalar masses: From the effective interactions in eq. (2), we can take the limit of integrating out the singlet scalar $S$ whose squared mass becomes $m^{2}_{S}\gg\kappa\kappa^{\prime}$. As a result, we obtain a renormalizable but $Z_{4}$ breaking effective interaction for two Higgs doublets from $\kappa,\kappa^{\prime}$ terms, at tree level as shown in Fig. 1, in the following form,
@@
Here, $\kappa$ is a $Z_{4}$ invariant dimensionful coupling, so it can be of order $m_{S}$. But, $\kappa^{\prime}$ is a $Z_{4}$ breaking term due to the VEV of the spurion field $\varphi$, so it is naturally suppressed as compared to $m_{S}$, that is, $\kappa^{\prime}=\lambda^{\prime}_{S\varphi}v_{\varphi}\ll m_{S}$. Therefore, the effective $\lambda_{5,{\rm eff}}$ term in eq. (4) is also suppressed as a consequence of the decoupled scalar.
@@
where $\lambda_{L}\equiv\lambda_{3}+\lambda_{4}$, with $\lambda_{3},\lambda_{4}$ being mixing quartic couplings in the scalar potential, $\lambda_{3}|H_{1}|^{2}|H_{2}|^{2}+\lambda_{4}(H^{\dagger}_{1}H_{2})(H^{\dagger}_{2}H_{1})$, and $m^{2}_{2}$ is the squared mass parameter for the inert doublet $H_{2}$. Henceforth, we choose $\lambda_{5,{\rm eff}}>0$ without loss of generality such that $A_{0}$ is a dark matter candidate.
In the decoupling limit of the singlet scalar $S$, the small effective $\lambda_{5}$ term gives rise to a mass splitting between two real neutral scalars of $H_{2}$ is suppressed as follows, $\Delta\equiv m_{H_{0}}-m_{A_{0}}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{2m_{0}},$ with $m^{2}_{0}\equiv(m^{2}_{H_{0}}+m^{2}_{A_{0}})/2$. The lighter neutral component of the inert doublet $H_{2}$ is a dark matter candidate avoiding the strong bound on DM-nucleon scattering with $Z$ boson exchange, as far as $|\Delta|\gtrsim 100\,{\rm keV}$, which sets a lower bound on $\lambda_{5,{\rm eff}}$ as $\lambda_{5,{\rm eff}}\gtrsim 3.3\times 10^{-7}(m_{0}/100\,{\rm GeV})(|\Delta|/100\,{\rm keV})$. Then, for $\kappa\sim m_{S}$, we need $\kappa^{\prime}/m_{S}\gtrsim 8\times 10^{-8}$ for $m_{0}\sim 100\,{\rm GeV}$ and $|\Delta|\gtrsim 100\,{\rm keV}$.
@@
Decoupled scalar and small neutrino masses: In the decoupling limit of the singlet scalar $S$, from the combination of the Yukawa interactions for right-handed neutrinos and a small effective $\lambda_{5,{\rm eff}}$ term in eq. (3), we obtain the one-loop corrections for neutrino masses from the Feynman diagram shown in Fig. 2 as
$({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}}\sum_{k}\frac{y_{N,ik}y_{N,jk}M_{N,k}}{m^{2}_{0}-M^{2}_{N,k}}\bigg{[}1-\frac{M^{2}_{N,k}}{m^{2}_{0}-M^{2}_{N,k}}\ln\frac{m^{2}_{0}}{M^{2}_{N,k}}\bigg{]}.$ (8)
The above results match with those due to a nonzero small $\lambda_{5}$ in the case without a $Z_{4}$ symmetry in Ref. Ma:2006km . Then, when the RH neutrinos are heavier than the inert doublet Higgs, namely, $M_{N,k}\gg m_{0}$, the above expression with eq. (4) gets simplified to
@@
Figure 3: Correlations for inert doublet scalar mass $m_{0}$ vs RH neutrino mass $M_{N}$ for varying $\lambda_{5,{\rm eff}}$ (left plot); neutrino mass $m_{\nu}$ vs singlet scalar mass $m_{S}$ for varying $M_{N}$ (right plot).
@@
We also comment on the case with $m_{0}\gg M_{N,k}$. In this case, the lightest RH neutrino is a dark matter candidate Ibarra:2016dlb whereas the scalars of the inert doublet decay into a pair of the SM lepton and the RH neutrino through the Yukawa interactions. In this limit, eq. (8) becomes approximated to $({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}m^{2}_{0}}\sum_{k}y_{N,ik}y_{N,jk}M_{N,k}.$ For instance, for $m_{\nu}\lesssim 0.1\,{\rm eV}$, $y_{N,ij}={\cal O}(1)$, $m_{0}\sim 100\,{\rm GeV}$ and $M_{N,k}\sim 10\,{\rm MeV}$, we can set the upper bound, $\lambda_{5,{\rm eff}}\lesssim 8\times 10^{-10}$ or $|\Delta|\lesssim 0.2\,{\rm keV}$, but there is no more lower bound on $\lambda_{5,{\rm eff}}$ because the scalars in the inert doublet decay. A small effective $\lambda_{5,{\rm eff}}$ is achievable if $m_{S}\gtrsim 10^{9}\,{\rm GeV}(\lambda^{\prime}_{S\varphi}/\lambda_{N,kk})$ for $v_{\varphi}=10\,{\rm MeV}/\lambda_{N,kk}$ for $\kappa\sim m_{S}$.
@@
In the left plot of Fig. 3, we show the parameter space for inert doublet scalar mass $m_{0}$ vs RH neutrino mass $M_{N}\equiv M_{N,1}$ satisfying $m_{\nu}=0.1\,{\rm eV}$. The color code indicates a scan over $\lambda_{5,{\rm eff}}$ between $10^{-8}$ and $10^{-4}$, showing that $\lambda_{5,{\rm eff}}\gtrsim 10^{-6}$ and $M_{N}\gtrsim 2\times 10^{7}\,{\rm GeV}$ are necessary for $m_{0}\gtrsim 100\,{\rm GeV}$. In the right plot of Fig. 3, we decode the correlation between neutrino mass $m_{\nu}$ and singlet scalar mass $m_{S}$ for varying $M_{N}$ between $10^{4}\,{\rm GeV}$ and $10^{9}\,{\rm GeV}$. In both plots in Fig. 3, we imposed $|\Delta|>100\,{\rm keV}$ and took into account the perturbativity, $|\lambda_{5,{\rm eff}}|<1$. Then, we need the singlet scalar mass $m_{S}$ to be of order $10^{14}-10^{16}\,{\rm GeV}$ for $m_{\nu}=0.01-0.1\,{\rm eV}$, which might hint at a connection of neutrino mass generation to the unification. If we choose smaller Yukawa couplings $y_{N}$, the diagonal boundary in the left plot of Fig. 3 is shifted to the left, allowing for smaller $M_{N}$ for a fixed $m_{0}$, and the strip in the right plot of Fig. 3 is shifted to the bottom, making $m_{S}$ smaller for a fixed $m_{\nu}$.

### v2

@@ Abstract
We propose a unified solution with $Z_{4}$ discrete symmetry for small neutrino masses and stability of dark matter. The Standard Model is extended with an inert doublet scalar, a dark singlet scalar, a spurion scalar and right-handed neutrinos, which all transform nontrivially under $Z_{4}$. After the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion, much below the mass scale of the dark singlet scalar, a small lepton number violating coupling for the inert doublet is generated at tree level, so small neutrino masses are obtained at one-loops. We discuss the important roles of the $Z_{4}$ symmetry for neutrino masses, dark matter physics and thermal leptogenesis.
I Introduction
@@
Neutrino oscillations ParticleDataGroup:2022pth indicate neutrinos masses of $m_{\nu}\sim 0.05\,{\rm eV}$. Moreover, the bounds from neutrinoless double beta decay experiments KATRIN:2021uub and Cosmic Microwave Background anisotropies and Baryon Acoustic Oscillations Planck:2018vyg are $m_{{\bar{\nu}}_{e}}<0.8\,{\rm eV}$ and $\sum_{i}m_{\nu_{i}}<0.12\,{\rm eV}$ at 95% CL, respectively. In order to account for small neutrino masses, it is possible to extend the Standard Model (SM) by the dimension-5 operators for neutrino masses with a cutoff scale beyond the electroweak scale or light sterile neutrinos coupled to neutrinos much below the electroweak scale Cai:2017jrq .
@@
Small neutrino masses can be obtained by the mixing between active neutrinos and heavy Majorana neutrinos Minkowski:1977sc ; Yanagida:1979as ; Gell-Mann:1979vob ; Mohapatra:1979ia . Alternatively, small neutrino masses can be generated by loop corrections, for instance, due to right-handed (RH) neutrinos and an extra doublet scalar Ma:2006km , the latter of which serves as a dark matter candidate Barbieri:2006dq ; LopezHonorez:2006gr . However, there is still a need of very small parameters in obtaining small neutrino masses in similar models Cai:2017jrq .
@@
In this article, we propose a novel model with a $Z_{4}$ discrete symmetry for the common origin of small neutrino masses and stability of dark matter. In our model, there are one inert doublet scalar, three right-handed neutrinos, a real spurion scalar and a singlet complex scalar field, all transforming nontrivially under $Z_{4}$. The lepton number is respected in the renormalizable interactions due to $Z_{4}$. Then, after the $Z_{4}$ symmetry is broken to $Z_{2}$ due to the VEV of the spurion field, the lepton number is broken, and the lightest neutral dark scalar in the model becomes a dark matter candidate. The spurion scalar can be regarded as a pseudo-Goldstone boson coming from the breaking of a global $U(1)$ symmetry into $Z_{4}$ at a higher scale, and $Z_{4}$ can be further down to $Z_{2}$ at a low scale.
@@
Taking the decoupling limit for making the singlet scalar field $S$ heavy, we show how a small mass splitting between the neutral components of the inert doublet arises, in turn giving rise to small neutrino masses at one-loops. We discuss the roles of the new singlet scalars and constraints for neutrino masses, dark matter and thermal leptogenesis.
II Discrete symmetry and dark matter scalars
@@
We introduce a complex singlet scalar $S$ and one extra electroweak doublet scalar, $H_{2}$, other than the SM Higgs doublet $H_{1}$, and three right-handed (RH) neutrinos, $N_{R,i}$, with $i=1,2,3$, and a spurion real scalar field $\varphi$, beyond the SM. We impose a $Z_{4}$ symmetry for the extra fields such that $\phi_{i}\to e^{iQ_{X}\pi/2}\phi_{i}$ with $\phi_{i}=S,\varphi,H_{1},H_{2},N_{R}$ and $Q_{X}(\phi_{i})=(+1,+2,0,+1,+1)$. But, the SM fermions are neutral under $Z_{4}$ so the model is different from the flavor-dependent $Z_{4}$ symmetry introduced in Ref. Babu:2013pma . The $Z_{4}$ symmetry can be embedded into a global $U(1)$ Peccei-Quinn symmetry which is anomalous in the presence of the couplings to the SM quarks or extra quarks, but we don’t pursue this possibility in detail in the current work. As the subgroup of the $Z_{4}$ symmetry, we can identify $Z_{2}$ parities as $Z_{2}(\phi_{i})=(-,+,+,-,-)$. The spurion field $\varphi$ transforms by $\varphi\to-\varphi$ under the $Z_{4}$ symmetry but it is invariant under the $Z_{2}$ subgroup.
@@
with ${\widetilde{H}}_{2}=i\sigma_{2}H^{*}_{2}$. It is remarkable that $\frac{1}{2}\lambda_{5}(H^{\dagger}_{1}H_{2})^{2}$ and its hermitian conjugate, vanish due to the $Z_{4}$ symmetry, and there is no bare mass terms for right-handed neutrinos due to the lepton number conservation. There are extra $Z_{4}$ invariant interaction terms in the potential such as $\mu\varphi S^{2},\frac{1}{2}\lambda_{H_{i}\varphi}\varphi^{2}|H_{i}|^{2}$($i=1,2$), etc, as shown in the appendix, but it is sufficient to focus on eq. (1) for our discussion below.
@@
where $\kappa^{\prime}\equiv\lambda^{\prime}_{S\varphi}v_{\varphi}$ is the $Z_{4}$ breaking effective interaction and $M_{N,i}\equiv\lambda_{N,ii}v_{\varphi}$ are the masses for RH neutrinos. Here, we assumed the basis where $\lambda_{N,ij}$ is diagonalized. After electroweak symmetry breaking with $\langle H_{1}\rangle=(0,v_{H})^{T}/\sqrt{2}\neq 0$, there appear mass mixings between dark matter scalars, namely, $S$ and the neutral components of $H_{2}$, so nonzero neutrino masses can be generated at one-loops, which will be discussed later. When the SM Higgs and the spurion field $\varphi$ are expanded around the VEVs as $H_{1}=(0,v_{H}+h)^{T}/\sqrt{2}$ and $\varphi=v_{\varphi}+\rho$, the dark Higgs $\rho$ can mix with the SM Higgs $h$, so the mass eigenstates become $h_{1}=c_{\alpha}\,h+s_{\alpha}\,\rho$ and $h_{2}=-s_{\alpha}\,h+c_{\alpha}\,\rho$, with $c_{\alpha}\equiv\cos\alpha,s_{\alpha}\equiv\sin\alpha$.
We remark that $S$ and $H_{2}$ do not get VEVs, leaving a $Z_{2}$ symmetry unbroken. $S$, $H_{2}$ and $N_{R,i}$ are odd under the remaining $Z_{2}$, which ensures the stability of the lightest neutral $Z_{2}$-odd particle among them as a dark matter candidate. For small neutrino masses without introducing small parameters, we consider the case where the singlet scalar $S$ is decoupled, so either the lighter neutral component of $H_{2}$ or the lightest RH neutrino can be a dark matter candidate.
III Decoupled scalar and inert doublet scalar masses
From the effective interactions in eq. (2), we can take the limit of integrating out the singlet scalar $S$ whose squared mass becomes $m^{2}_{S}\gg\kappa\kappa^{\prime}$. As a result, we obtain a renormalizable but $Z_{4}$ breaking effective interaction for two Higgs doublets from $\kappa,\kappa^{\prime}$ terms, at tree level as shown in Fig. 1, in the following form,
@@ III Decoupled scalar and inert doublet scalar masses
Here, $\kappa$ is a $Z_{4}$ invariant dimensionful coupling, so it can be of order $m_{S}$. But, $\kappa^{\prime}$ is a $Z_{4}$ breaking coupling due to the VEV of the spurion field $\varphi$, so it is naturally suppressed as compared to $m_{S}$, that is, $\kappa^{\prime}=\lambda^{\prime}_{S\varphi}v_{\varphi}\ll m_{S}$. Therefore, the effective $\lambda_{5,{\rm eff}}$ term in eq. (4) is also suppressed as a consequence of the decoupled scalar.
Figure 1: Feynman diagrams contributing to $\lambda_{5,{\rm eff}}$.
@@
where $\lambda_{L}\equiv\lambda_{3}+\lambda_{4}$, with $\lambda_{3},\lambda_{4}$ being mixing quartic couplings in the scalar potential, $\lambda_{3}|H_{1}|^{2}|H_{2}|^{2}+\lambda_{4}(H^{\dagger}_{1}H_{2})(H^{\dagger}_{2}H_{1})$, and $m^{2}_{2}$ is the squared mass parameter for the inert doublet $H_{2}$. Henceforth, we choose $\lambda_{5,{\rm eff}}>0$ without loss of generality such that $A_{0}$ is a dark matter candidate. The most general results for the mass eigenvalues and the mixing angles with the non-decoupled singlet scalar $S$ are shown in eqs. (39), (50) and (51), in the appendix.
In the decoupling limit of the singlet scalar $S$, the small effective $\lambda_{5}$ term gives rise to a mass splitting between two real neutral scalars of $H_{2}$ is suppressed as $\Delta\equiv m_{H_{0}}-m_{A_{0}}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{2m_{0}},$ with $m^{2}_{0}\equiv(m^{2}_{H_{0}}+m^{2}_{A_{0}})/2$. So, we can avoid the strong bound on DM-nucleon scattering with $Z$ boson exchange, as far as $|\Delta|\gtrsim 100\,{\rm keV}$, which sets a lower bound on $\lambda_{5,{\rm eff}}$ as $\lambda_{5,{\rm eff}}\gtrsim 3.3\times 10^{-7}(m_{0}/100\,{\rm GeV})(|\Delta|/100\,{\rm keV})$. Then, for $\kappa\sim m_{S}$, we need $\kappa^{\prime}/m_{S}\gtrsim 8\times 10^{-8}$ for $m_{0}\sim 100\,{\rm GeV}$ and $|\Delta|\gtrsim 100\,{\rm keV}$.
@@
IV Decoupled scalar and small neutrino masses
In the decoupling limit of the singlet scalar $S$, from the combination of the Yukawa interactions for right-handed neutrinos and a small effective $\lambda_{5,{\rm eff}}$ term in eq. (3), we obtain the one-loop corrections for neutrino masses from the Feynman diagram shown in Fig. 2 as
$({\cal M}_{\nu})_{ij}$ $\simeq$ $\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}}\sum_{k}\frac{y_{N,ik}y_{N,jk}M_{N,k}}{m^{2}_{0}-M^{2}_{N,k}}$ (8)
×$\times\bigg{[}1-\frac{M^{2}_{N,k}}{m^{2}_{0}-M^{2}_{N,k}}\ln\frac{m^{2}_{0}}{M^{2}_{N,k}}\bigg{]}.$
The above results match with those due to a nonzero small $\lambda_{5}$ in the case without a $Z_{4}$ symmetry in Ref. Ma:2006km , and the same results can be obtained from the decoupling limit of the general results in eq. (53) in the appendix. Then, when the RH neutrinos are heavier than the inert doublet Higgs, namely, $M_{N,k}\gg m_{0}$, the above expression with eq. (4) gets simplified to
@@ IV Decoupled scalar and small neutrino masses
In the case with $m_{0}\gg M_{N,k}$, the lightest RH neutrino is a dark matter candidate Ibarra:2016dlb whereas the scalars of the inert doublet decay into a pair of the SM lepton and the RH neutrino through the Yukawa interactions. In this limit, eq. (8) becomes approximated to $({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}m^{2}_{0}}\sum_{k}y_{N,ik}y_{N,jk}M_{N,k}.$ For instance, for $m_{\nu}\lesssim 0.1\,{\rm eV}$, $y_{N,ij}={\cal O}(1)$, $m_{0}\sim 100\,{\rm GeV}$ and $M_{N,k}\sim 10\,{\rm MeV}$, we can set the upper bound, $\lambda_{5,{\rm eff}}\lesssim 8\times 10^{-10}$ or $|\Delta|\lesssim 0.2\,{\rm keV}$. But, there is no more lower bound on $\lambda_{5,{\rm eff}}$ because the scalars in the inert doublet decay. A small effective $\lambda_{5,{\rm eff}}$ is achievable if $m_{S}\gtrsim 10^{9}\,{\rm GeV}(\lambda^{\prime}_{S\varphi}/\lambda_{N,kk})$ for $v_{\varphi}=10\,{\rm MeV}/\lambda_{N,kk}$ for $\kappa\sim m_{S}$.
@@
In Fig. 3, we show the parameter space for inert doublet scalar mass $m_{0}$ vs RH neutrino mass $M_{N}\equiv M_{N,1}$ satisfying $m_{\nu}=0.1\,{\rm eV}$. The color code indicates a scan over $\lambda_{5,{\rm eff}}$ between $10^{-8}$ and $10^{-4}$, showing that $\lambda_{5,{\rm eff}}\gtrsim 10^{-6}$ and $M_{N}\gtrsim 2\times 10^{7}\,{\rm GeV}$ are necessary for $m_{0}\gtrsim 100\,{\rm GeV}$. In Fig. 4, we decode the correlation between neutrino mass $m_{\nu}$ and singlet scalar mass $m_{S}$ for varying $M_{N}$ between $10^{4}\,{\rm GeV}$ and $10^{9}\,{\rm GeV}$. In both Figs. 3 and 4, we imposed $|\Delta|>100\,{\rm keV}$ and took into account the perturbativity, $|\lambda_{5,{\rm eff}}|<1$. Then, we need the singlet scalar mass $m_{S}$ to be of order $10^{14}-10^{16}\,{\rm GeV}$ for $m_{\nu}=0.01-0.1\,{\rm eV}$, which might hint at a connection of neutrino mass generation to the unification. If we choose smaller Yukawa couplings $y_{N}$, the diagonal boundary in Fig. 3 is shifted to the left, allowing for smaller $M_{N}$ for a fixed $m_{0}$, and the strip in Fig. 4 is shifted to the bottom, making $m_{S}$ smaller for a fixed $m_{\nu}$.
Figure 3: Inert doublet scalar mass $m_{0}$ vs RH neutrino mass $M_{N}$ for varying $\lambda_{5,{\rm eff}}$.
Figure 4: Neutrino mass $m_{\nu}$ vs singlet scalar mass $m_{S}$ for varying $M_{N}$
V Dark matter direct detection